OpenStudy (anonymous):

What is the solution to the rational equation 4/3 + 3/ x+5 = 21/ 3x+15 ? someone please help

OpenStudy (pfenn1):

\[\left( 4 \over 3 \right)+\left( 3 \over x+5 \right)=\left( 21 \over 3x+5 \right)\] what would be the common denominators for all three terms in the equation.

OpenStudy (anonymous):

there isnt one

OpenStudy (pfenn1):

Sorry. I made a typo on the equation.\[\left( 4 \over 3 \right)+\left( 3 \over x+5 \right)=\left( 21\over 3x +15 \right)\]

OpenStudy (pfenn1):

Common denominator?

OpenStudy (anonymous):

3

OpenStudy (pfenn1):

Close. Hint: 3x+15 can be factored.

OpenStudy (anonymous):

x+5

OpenStudy (pfenn1):

Close. If you have an equation\[\left( 1\over a \right) + \left( 1 \over b \right)=\left( 1 \over ab \right)\]what would be the common denominator

OpenStudy (anonymous):

ab

OpenStudy (pfenn1):

Right, you have the same thing here. a is like 3 and b is like x+5. So what is the common denominator of your equation?

OpenStudy (anonymous):

3x+5

OpenStudy (pfenn1):

Close. It would be 3(x+5)=3x+15

OpenStudy (anonymous):

ok

OpenStudy (pfenn1):

So you need to get each term to be over the common denominator. We multiply the first term by x+5/x+5 to get\[\left( 4 \over 3 \right)\left( x+5 \over x+5 \right)=?\]

OpenStudy (anonymous):

4(x+5) / 3(x+5)

OpenStudy (pfenn1):

So what would you multiply the second term by to get the common denominator?

OpenStudy (anonymous):

4?

OpenStudy (pfenn1):

You want to end up with a term that has the common denominator on the bottom 3(x+5) so you would multiply by 3/3 \[\left( 3 \over x+5 \right)\left( 3 \over 3 \right)=?\]

OpenStudy (anonymous):

9/ 3x+5

OpenStudy (pfenn1):

9/3(x+5)so now our equation looks like\[\left( 4(x+5)+9 \over 3 (x+5) \right)=\left( 21 \over 3 (x+5) \right)\] Can you rearrange and solve for x?